how many gallons 65%antifreeze and how many 20% antifreeze should be mixed up to obtain 50 gallons of 56% mixture of antifreeze?
Answer by mananth(16946)

percent ---------------- quantity
Antifreeze I 65----------------x gallons
Antifreeze II 20 ----------------50-x gallons
Total 56.00% ----------------50 gallons

65x+20(50-x)=56 *50
65x+1000-20x=2800
65x-20x=2800-1000
45x=1800
/45
x=40 gallons 65.00%
10 gallons 20.00%